OpenStudy (anonymous):

What are the number of solutions to the system x2 - 2y2 = 12 and 2x2 + y2 = 4? (NOTE: JUST THE NUMBER OF SOLUTIONS.)

OpenStudy (welshfella):

if we multiply second equation by 2 we get 4x^2+ 2y^2 = 8 adding this to equation 1 gives us an equation in x only which we can solve for x so what do you think ?

OpenStudy (anonymous):

1 Solution only?

OpenStudy (anonymous):

Oh ok, Im thinking no solutions

OpenStudy (welshfella):

hmm let me think a bit more about this one..

OpenStudy (welshfella):

we get 5x^2 = 20 x = +2 , -2

OpenStudy (welshfella):

plugging 2 or -2 into equation 1 gives 4 - 2y2 = 12 2y^2 = -8 y^2 = -1 - no real values of y so there is no real value of y corresponding to 2 or -2 yes you are right - no real solution

OpenStudy (anonymous):

Thank You!

OpenStudy (welshfella):

yw